Please help with this question if possible

Solve x^2 - 7x - 3 = 0 for x using any method you prefer
john

Answer by richwmiller(17219)

x^2 - 7x - 3 = 0
complete the square
(x-7/2)^2=61/4
I leave the rest to you.